# Text-to-Vision Generation with Scene Graph Programming

### v2

@@ 5 Application 1: Self-Improving Models
Figure 4: Results for Application 1: Self-Improving Models. Average VQA score of Stable Diffusion v1-5 fine-tuned on different data across 1K Generate Any Scene image/video evaluation set and GenAI-Bench image/video benchmark [34].
@@
In this section, we explore how Generate Any Scene facilitates a self-improvement framework for model generation capabilities. By programmatically generating scalable compositional captions from scene graphs, Generate Any Scene expands the textual and visual space, allowing for a diversity of synthetic images that extend beyond real-world scenes. Our goal is to utilize these richly varied synthetic images to further boost model performance.
Iterative self-improving framework. Inspired by DreamSync [62], we designed an iterative self-improving framework using Generate Any Scene with Stable Diffusion v1-5 as the baseline model. With VQA Score, which shows strong correlation with human evaluations on compositional images [39], we guide the model’s improvement throughout the process.
@@
Specifically, Generate Any Scene generates 3 $\times$ 10K captions across three epochs. For each caption, Stable Diffusion v1-5 generates 8 images, and the image with the highest VQA Score is selected. From each set of 10K optimal images, we then select the top 25% (2.5k image-caption pairs) as the training data for each epoch. In subsequent epochs, we use the fine-tuned model from the prior iteration to generate new images. We employ LoRA [21] for parameter-efficient fine-tuning. Additional details are available in Appendix E.
@@
To evaluate the effectiveness of self-improvement using synthetic data generated by Generate Any Scene, we conduct comparative experiments with the CC3M dataset, which comprises high-quality and diverse real-world image-caption pairs [60]. We randomly sample 3 $\times$ 10K captions from CC3M, applying the same top-score selection strategy for iterative fine-tuning of Stable Diffusion v1-5. Additionally, we include a baseline using random-sample fine-tuning strategy to validate the advantage of our highest-scoring selection-based strategy.
Results. We evaluate our self-improving pipeline on Text-to-Vision generation benchmarks, including GenAI Bench [34]. For the Text-to-Video generation task, we use Text2Video-Zero as the baseline model, substituting its backbone with the original Stable Diffusion v1-5 and our fine-tuned Stable Diffusion v1-5 models.
@@
Our results show that fine-tuning with Generate Any Scene-generated synthetic data consistently outperforms CC3M-based fine-tuning across Text-to-Vision generation tasks (Figure 4), achieving the highest gains with our highest-scoring selection strategy. This highlights Generate Any Scene’s scalability and compositional diversity, enabling models to effectively capture complex scene structures. Additional results are in Appendix F.
Takeaway for application 1 Iterative self-improving Text-to-Vision generation models with compositional and diverse synthetic captions can surpass fine-tuning with real-world image-caption data. Potential reason: The compositional, synthetic captions generated by Generate Any Scene exhibit greater diversity than real-world data.
6 Application 2: Distilling limitations
@@
Although self-improving with Generate Any Scene-generated data shows clear advantages over high-quality real-world datasets, its efficiency remains inherently constrained by the limitations of the model’s own generation ability. To address this, we leverage the taxonomy and programmatic generation capabilities within Generate Any Scene to identify specific strengths of proprietary models (DaLL-E 3), and to distill these capabilities into open-source models. More details are in Appendix F.
6.1 Fine-Grained Analysis of DaLL-E 3’s Exceptional Performance
Figure 5: Examples for Application 2: Distilling limitations. Examples of images generated by DaLL-E 3, the original Stable Diffusion v1-5, and the fine-tuned versions. The left four captions demonstrate fine-tuning with multi-object captions generated by Generate Any Scene for better compositionality, while the right two columns focus on understanding hard concepts.
@@ 6.1 Fine-Grained Analysis of DaLL-E 3’s Exceptional Performance
As shown in Figure 3, DaLL-E 3 achieves TIFA Score 1.5 to 2 times higher than those of other models. When we compare TIFA Score across varying numbers of elements (objects, relations, and attributes per caption) in Figure 6(b), DaLL-E 3 counterintuitively maintains consistent performance regardless of element count. The performance of other models declines as the element count increases, which aligns with expected compositional challenges. We suspect these differences are primarily due to DaLL-E 3’s advanced capabilities in compositionality and understanding hard concepts, which ensures high faithfulness across diverse combinations of element types and counts.
6.2 Distilling compositionality from DaLL-E 3
Observations. We find that DaLL-E 3 tends to produce straightforward combinations of multiple objects (Figure 5). In contrast, open-source models like Stable Diffusion v1-5 often omit some objects from the captions, even though they are capable of generating each object individually.
This difference suggests that DaLL-E 3 may be trained on datasets emphasizing multi-object presence without rigorous attention to image layout or object interaction. Such training likely underpins DaLL-E 3’s stronger performance on metrics like TIFA Score and VQA Score, prioritizing object inclusion over detailed compositional arrangement.
Finetuning. To encourage Stable Diffusion v1-5 to learn compositional abilities similar to those of DaLL-E 3., we select a set of 778 images generated by DaLL-E 3, each containing multiple objects, and utilize this dataset to fine-tune Stable Diffusion v1-5. For the baseline, we randomly sampled an equivalent set of DaLL-E 3-generated images paired with generated captions from Generate Any Scene.
Results. To evaluate compositional improvements, we generate 1K multi-object captions. Figure 6(b) shows a 10% TIFA Score increase after fine-tuning, random fine-tuning by an average of 3%. These results indicate enhanced compositional abilities in handling complex generation tasks.
We analyze images generated by Stable Diffusion v1-5 before and after fine-tuning on high-complexity image-caption pairs (Figure 5). It is surprising to see that, with only 1K LoRA fine-tuning steps, Stable Diffusion v1-5 effectively learn DaLL-E 3 ’s capability to arrange and compose multiple objects within a single image,. This fine-tuning strategy notably enhances alignment between generated images and their given captions.
On a broader set of 10K Generate Any Scene-generated captions, the fine-tuned model consistently outperformed the randomly fine-tuned model (Figure 6(a)), confirming the generalizability and superiority of targeted fine-tuning for improving model performance.
@@ 6.2 Distilling compositionality from DaLL-E 3
(a) Distilling compositionality from DaLL-E 3: Model results on TIFA vs. total element numbers in captions in 10K general Generate Any Scene captions.
@@
(c) Learning hard concepts from DALL-E 3: Models’ average TIFA Score performance over captions and hard concepts in 1K hard concepts Generate Any Scene captions.
@@
Figure 6: Results for Application 2: Distilling limitations. The left two figures show the results for Distilling compositionality from DALL-E 3, while the rightmost figure shows the results for Learning hard concepts from DALL-E 3.
6.3 Learning hard concepts from DaLL-E 3
Observation. Figure 5 shows that is capable not only of handling multi-object generation but also of understanding and generating rare and hard concepts, such as a specific species of flower. We attribute this to its training with proprietary real-world data.
Finetuning. Using the taxonomy of Generate Any Scene, we compute model performance on each conceptby averaging scores across captions containing that concept.Accumulating results through the taxonomy, we identify the 100 concepts where Stable Diffusion v1-5 shows the largest performance gap relative to DaLL-E 3. For fine-tuning, we generate 778 captions incorporating these concepts with others, using DaLL-E 3 to produce corresponding images. As a baseline, we randomly select 778 Generate Any Scene-generated captions for fine-tuning and compare these with the original Stable Diffusion v1-5 model.
@@ 6.3 Learning hard concepts from DaLL-E 3
Figure 7: Results for Application 3: Generated content detector. Comparison of detection performance across different data scales using $D^{3}$ alone versus the combined $D^{3}$ + Generate Any Scene training set in cross-model and cross-dataset scenarios.
Results. The results in Figure 6(c) show that our targeted fine-tuning led to improved model performance, reflected in higher average scores across captions and increased scores for each challenging concept.
Takeaway for application 2 Targeted fine-tuning can distill proprietary model strengths, effectively bridging gaps in compositionality and concept handling for open-source models. Potential Reason: Generate Any Scene facilitates fine-grained analysis to identify specific performance gaps, enabling targeted data selection to distill limitations.
7 Application 3: Generated content detector
Advances in Text-to-Vision generation underscore the need for effective content moderation [50]. Major challenges include the lack of high-quality and diverse datasets and the difficulty of generalizing detection across models Text-to-Vision generation [68, 29]. Generate Any Scene addresses these issues by enabling scalable, programmatic generation of compositional captions, increasing the diversity and volume of synthetic data. This approach enhances existing datasets by compensating for their limited scope-from realistic to imaginative-and variability.
To demonstrate Generate Any Scene’s effectiveness in training generated content detectors, we used the $D^{3}$ dataset [3] as a baseline. We sampled 5k captioned real and SDv1.4-generated image pairs from $D^{3}$ and generated 5k additional images with Generate Any Scene captions. We trained a ViT–T [73] model with a single-layer linear classifier, varying dataset sizes with $N$ real and $N$ synthetic images. For synthetic data, we compared $N$ samples solely from $D^{3}$ with a mixed set of $N/2$ from Generate Any Scene and $N/2$ from $D^{3}$, keeping the same training size.
We evaluate the detector’s generalization on the GenImage [85] validation set and images generated using Generate Any Scene captions. Figure 7 demonstrates that combining Generate Any Scene-generated images with real-world captioned images consistently enhances detection performance, particularly across cross-model scenarios and diverse visual scenes. More details are in Appendix G.
Takeaway for application 3 Compositional synthetic captions robustify generated content detectors. Potential reason: Generate Any Scene can generate more diverse captions to complement real-world image-caption training data by enriching compositional variety and imaginative scope.
8 Limitation
Programmatically generated prompts can be unrealistic and biased.
Linguistic diversity of programmatic prompts is limited.

### v3

@@ 3 Self-Improving models with synthetic captions
Figure 2: Results for Self-Improving Models. Average VQA score of SDv1.5 fine-tuned on different data across 1K Generate Any Scene image/video evaluation set and GenAI-Bench image/video benchmark (Li et al., 2024a).
@@
With Generate Any Scene, we develop a self-improvement framework to improve generative capabilities. By generating scalable compositional captions from scene graphs, Generate Any Scene expands the textual and visual space, allowing for a diversity of synthetic images that extend beyond real-world scenes. Our goal is to utilize these richly varied synthetic images to further boost model performance.
@@
Iterative self-improving framework. Inspired by DreamSync (Sun et al., 2023), we designed an iterative self-improving framework using Generate Any Scene with SDv1.5 as the baseline model. With VQA Score, which shows strong correlation with human evaluations on compositional images (Lin et al., 2024), we guide the model’s improvement throughout the process. Specifically, Generate Any Scene generates 3 $\times$ 10K captions across three epochs. For each caption, SDv1.5 generates 8 images, and the image with the highest VQA Score is selected. From each set of 10K optimal images, we then select the top 25% (2.5K image-caption pairs) as the training data for each epoch. In subsequent epochs, we use the fine-tuned model from the prior iteration to generate new images. We employ LoRA (Hu et al., 2021) for parameter-efficient fine-tuning.
@@
Baselines. We conduct comparative experiments with the CC3M dataset, which comprises high-quality and diverse real-world image-caption pairs (Sharma et al., 2018). We randomly sample 3 $\times$ 10K captions from CC3M, applying the same top-score selection strategy for iterative fine-tuning of SDv1.5. Additionally, we include a baseline using random-sample fine-tuning strategy to validate the advantage of our highest-scoring selection-based strategy. We evaluate our self-improving pipeline on Text-to-Vision generation benchmarks, including GenAI Bench (Li et al., 2024a). For the Text-to-Video generation task, we use Text2Video-Zero as the baseline model, substituting its backbone with the original SDv1.5 and our fine-tuned SDv1.5 models.
@@
Fine-tuning with our synthetic captions can surpass high-quality real-world image-caption data. Our results show that fine-tuning with Generate Any Scene-generated synthetic data consistently outperforms CC3M-based fine-tuning across Text-to-Vision generation tasks (Figure 2), achieving the highest gains with our highest-scoring selection strategy. This highlights Generate Any Scene’s scalability and compositional diversity, enabling models to effectively capture complex scene structures. Additional experiment settings and results are in Appendix C.
4 Distilling targeted capabilities
@@
Although self-improving with Generate Any Scene shows clear advantages over high-quality real-world datasets, its efficiency is inherently limited by the model’s own generation capabilities. To address this, we leverage the taxonomy and systematical generation capabilities within Generate Any Scene to identify specific strengths of proprietary models (DaLL-E 3), and distill these capabilities into open-source models. More details are in Appendix D.
@@
We evaluate multiple models using Generate Any Scene controllably generated captions and observe that DaLL-E 3 achieves TIFA Score 1.5 to 2 times higher than those of other models. As shown in Figure 4(a), when comparing TIFA Score across captions with varying numbers of elements (objects, relations, and attributes), DaLL-E 3 counterintuitively maintains consistent performance regardless of element count. The performance of other models declines as the element count increases, which aligns with expected compositional challenges. We suspect that these differences are primarily due to DaLL-E 3’s advanced capabilities in compositionality and understanding hard concepts, which ensures high faithfulness across diverse combinations of element types and counts.
Figure 3: Examples for Distilling Capabilities. Examples of images generated by DaLL-E 3, the original SDv1.5, and the fine-tuned versions. The left four captions demonstrate fine-tuning with multi-object captions generated by Generate Any Scene for better compositionality, while the right two columns focus on understanding hard concepts.
Distilling compositionality from DaLL-E 3. When analyzing model outputs from our synthetic captions, we find that DaLL-E 3 tends to produce straightforward combinations of multiple objects (Figure 3). In contrast, open-source models like SDv1.5 often omit objects from the captions, despite being capable of generating each one individually. This difference suggests that DaLL-E 3 may benefit from training data emphasizing multi-object presence, even without detailed layout or object interaction. Such training likely underpins DaLL-E 3’s stronger performance on metrics like TIFA Score and VQA Score that prioritize object inclusion. To effectively distill these compositional abilities into SDv1.5, we employ Generate Any Scene for targeted synthesis of 778 multi-object captions, paired with images generated by DaLL-E 3, for finetuning SDv1.5.
Distilling hard concepts understanding from DaLL-E 3. Figure 3 shows that DaLL-E 3 is capable not only of handling multi-object generation but also of understanding and generating rare and hard concepts, such as a specific species of flower. We attribute this to its training with proprietary real-world data. Using the taxonomy of Generate Any Scene, we compute model performance on each concept by averaging generation scores across captions containing that concept. Accumulating results through the taxonomy, we identify the 100 concepts where SDv1.5 shows the largest performance gap relative to DaLL-E 3. For distilling, we generate 778 captions incorporating these hard concepts with other elements, and use DaLL-E 3 to produce corresponding images.
@@
(a) Distilling compositionality from DaLL-E 3: Model results on TIFA vs. total element numbers in captions in 10K general Generate Any Scene captions. ("Best Open-Source Model" refers to Flux.1-schnell)
@@
(c) Distilling hard concepts understanding from DALL-E 3: Models’ average TIFA Score performance over captions and hard concepts in 1K hard concepts Generate Any Scene captions.
@@
Figure 4: Results for Distilling Capabilities. The left two figures show the results for Distilling compositionality, while the rightmost figure shows the results for Distilling hard concepts understanding from DALL-E 3.
Baselines. For the baseline, we randomly synthesize 778 captions using Generate Any Scene paired with DaLL-E 3-generated images to fine-tune the model. To evaluate model improvements, we generate another 1K multi-object captions and 1K hard-concept captions separately.
Targeted caption synthesis via Generate Any Scene enables effective distillation of compositional abilities and hard concept understanding. We analyze images generated by SDv1.5 before and after fine-tuning on high-complexity captions (Figure 3). Surprisingly, with fewer than 1K LoRA fine-tuning steps, SDv1.5 effectively learns DaLL-E 3 ’s capability to arrange and compose multiple objects within a single image. Quantitatively, Figure 4(b) shows a 10% improvement in TIFA Score after targeted fine-tuning, surpassing the performance of the randomly fine-tuned model. On a broader set of 10K Generate Any Scene-generated captions, the targeted fine-tuned model consistently outperforms randomly fine-tuned and original counterparts across complex scenes (Figure 4(a)). These results confirm not only the effectiveness but also the scalability and efficiency of Generate Any Scene. Also, the results in Figure 4(c) show that our targeted fine-tuning with hard concepts leads to improved model performance, reflected in higher average scores across captions and increased scores for each challenging concept.
5 Reinforcement learning with a synthetic reward function
Reinforcement Learning with Human Feedback (RLHF) has become an increasingly popular fine-tuning strategy in text-to-image generation (Gong et al., 2025; Jiang et al., 2025; Wang et al., 2025). However, defining an effective reward model that accurately captures semantic alignment for text-to-image generation remains an open challenge. Existing reward models like CLIP offer only coarse-grained image-text similarity signals, which fall short in assessing compositional correctness and lack interpretability. Alternative approaches have explored using visual question answering (VQA) as a proxy for evaluating semantic alignment, aiming for finer-grained assessments, yet require either labor-intensive datasets with dense annotations or large volumes of contextually relevant questions via advanced LLMs. Leveraging its structured scene graph synthesis capabilities, Generate Any Scene offers a scalable alternative by producing exhaustive semantic queries with negligible overhead, enabling low-cost, compositional reward modeling (Sec 2.1).
Figure 5: Comparison of generated images. Our reward model enables image generation with better semantic alignment, realism, and visual quality than baselines.
Experiment setup. Building on this scene graph-based reward modeling strategy, we adopt Group Relative Policy Optimization (GRPO) as our reinforcement learning algorithm. We fine-tune the SimpleAR-0.5B-SFT model for one epoch using 10K captions generated by Generate Any Scene, each paired with their scene graph-derived QA sets. For reward evaluation, we use Qwen2.5-VL-3B, a lightweight open-source vision-language model, to answer these QA pairs given the model-generated images. The reward is computed as the accuracy across all questions. This fine-grained, scene graph-aligned reward provides precise feedback on compositional faithfulness. As a baseline, we compare against SimpleAR-0.5B-RL, trained with CLIP-based rewards on 11K captions from real world datasets for one epoch. We evaluate our scene graph-based reward model on three benchmarks: DPG-Bench (Hu et al., 2024a), GenEval (Ghosh et al., 2023), and GenAI-Bench (Li et al., 2024a). More details are in Appendix E.
Generate Any Scene rewards outperform CLIP. As shown in Table 2, our method outperforms both SFT and CLIP-RL models and achieves a significant improvement, demonstrating superior compositional faithfulness driven by explicit scene graph rewards. Importantly, this performance gain is directly enabled by the Generate Any Scene engine, which constructs explicit scene graphs to generate compositional captions. Generate Any Scene provides a structured and cognitively aligned visual representation, from which we derive exhaustive QA pairs with minimal additional cost. Combined with lightweight VLM judge, this approach offers a scalable, low-cost solution for semantic-level reward modeling.
Table 2: Evaluation on the DPG, GenEval and GenAI benchmark. GRPO training with our reward model outperforms both SFT baseline and CLIP-RL models. TO: two objects, P: position, CA: color attribute.
Method DPG-Bench GenEval GenAI-Bench
Global Relation Overall TO P CA Overall Basic Advanced All
SimpleAR-0.5B-SFT
85.02 86.59 78.48 0.73 0.22 0.23 0.53 0.74 0.60 0.66
SimpleAR-0.5B-RL (Clip)
86.64 88.51 79.66 0.82 0.26 0.38 0.59 0.75 0.60 0.67
SimpleAR-0.5B-RL (Ours)
88.46 90.13 80.50 0.81 0.31 0.38 0.61 0.75 0.61 0.68
6 Improving generated-content detection
Advances in Text-to-Vision generation underscore the need for effective content moderation (Pei et al., 2024). Major challenges include the lack of high-quality and diverse datasets and the difficulty of generalizing detection across models Text-to-Vision generation (Wang et al., 2024b; Kaur et al., 2024). Generate Any Scene addresses these issues by enabling scalable, systematical generation of compositional captions, increasing the diversity and volume of synthetic data. This approach enhances existing datasets by compensating for their limited scope-from realistic to imaginative-and variability.
Experiment setup. To demonstrate Generate Any Scene’s effectiveness in training generated content detectors, we used the $D^{3}$ dataset (Baraldi et al., 2024) as a baseline. We sampled 5K captioned real and SDv1.4-generated image pairs from $D^{3}$ and generated 5K additional images with Generate Any Scene captions. We trained a ViT–T (Wu et al., 2022) model with a single-layer linear classifier, and compared models trained with samples solely from $D^{3}$ against those trained with samples Generate Any Scene and $D^{3}$.
Generate Any Scene improves generated content detectors. We evaluate the detector’s generalization on the GenImage (Zhu et al., 2024) validation set and images generated using Generate Any Scene captions. Figure 6 demonstrates that combining Generate Any Scene-generated images with real-world captioned images consistently enhances detection performance, particularly across cross-model scenarios and diverse visual scenes. More details are in Appendix F.
@@ 6 Improving generated-content detection
Figure 6: Results for Application 4: Generated content detector. Comparison of detection performance across different data scales using $D^{3}$ alone versus the combined $D^{3}$ + Generate Any Scene training set in cross-model and cross-dataset scenarios.
7 Comprehensive evaluation with Generate Any Scene
Beyond showcasing Generate Any Scene in model training, we also show that Generate Any Scene is a valuable resource for comprehensive and compositional evaluation. Specifically, we synthesize 10K captions for text-to-image, 10K for text-to-video, and 1K for text-to-3D, covering diverse scene structures and content topics. We evaluate 12 text-to-image, 9 text-to-video, and 5 text-to-3D models. Evaluations combine Generate Any Scene synthetic scene graphs with existing metrics (e.g., CLIP Score (Chakrabarty et al., 2023), VQA Score (Lin et al., 2024), TIFA Score (Hu et al., 2023; Cho et al., 2023)) to assess semantic similarity, faithfulness, and human preference alignment. Our key findings include: (1) DiT-backbone text-to-image models align more closely with input captions than UNet-backbone models. (2) Text-to-video models struggle with balancing dynamics and consistency, while both text-to-video and text-to-3D models show notable gaps in human preference alignment. Except for aggregating quantitative results, we also leverage Generate Any Scene ’s controllable captioning to evaluate models on fine-grained factors: perplexity, scene complexity, commonsense reasoning, and content category variation for case study.
Overall, Generate Any Scene yields stable, human-aligned rankings across T2I/T2V/T2-3D. Through broad, controllable coverage of objects, attributes, relations, and categories, it serves as a compositional stress test that reliably exposes plausibility gaps, category brittleness, and long-tail concept failures in current models (see Appendix A).
8 Related work
Text-to-Vision generation models.
Text-to-Image generation advances are driven by diffusion models and LLMs. Some open-source models (Rombach et al., 2022b; Podell et al., 2023; Li et al., 2024b; Pernias et al., 2023; Rombach et al., 2022a; at StabilityAI, 2023) use UNet backbones to refine images iteratively. In parallel, Diffusion Transformers (DiTs) architectures(Esser et al., 2024; Chen et al., 2023a; 2024b; Labs, 2024) have emerged as a better alternative in capturing long-range dependencies and improving coherence. Proprietary models like DALL-E 3 (Betker et al., 2023) and Imagen 3 (Baldridge et al., 2024) still set the state-of-the-art. Based on Text-to-Image generation method, Text-to-Video generation models typically utilize time-aware architectures to ensure temporal coherence across frames (Guo et al., 2023b; Wang et al., 2024a; Khachatryan et al., 2023; Wang et al., 2023b; Wu et al., 2025; Chen et al., 2024a; Yang et al., 2024; Zheng et al., 2024). In Text-to-3D generation, recent proposed models (Wang et al., 2024c; Poole et al., 2022; Wang et al., 2023a; Metzer et al., 2023; Lin et al., 2023) integrate the diffusion models with Neural Radiance Fields (NeRF) rendering to generate diverse 3D objects. Recent studies (Wang et al., 2025; Jiang et al., 2025; Wu et al., 2024a; Chen et al., 2025) have also explored the integration of image generation into a unified multimodal language model (MLM) framework based on auto-regressive transformer architectures, demonstrating promising improvements in both performance and efficiency.
Synthetic captions for Text-to-Vision generation.
Captions for Text-to-Vision generation models vary greatly in diversity, complexity, and compositionality. This variation makes it challenging and costly to collect large-scale and diverse captions written by humans. Consequently, synthetic captions have been widely used for both training (Lian et al., 2023; Sun et al., 2023; Li et al., 2024c; Zhao et al., 2024; Sun et al., 2024; Park et al., 2021; Wen et al., 2023; Wu et al., 2024b) and evaluation purposes (Huang et al., 2023). For example, training methods like LLM-Grounded Diffusion (Lian et al., 2023) leverage LLM-generated captions to enhance the model’s understanding and alignment with human instruction. For evaluation, benchmarks such as T2I-CompBench (Huang et al., 2023) and T2V-CompBench (Sun et al., 2024) utilize benchmarks generated by LLMs. However, LLMs are hard to control and may introduce exhibit systematic bias. In this work, we propose a programmatic scene graph-based data engine that can generate infinitely diverse captions for improving Text-to-Vision generation models.
Finetuning techniques for Text-to-Vision generation.
To accommodate the diverse applications and personalization needs in text-to-vision models, numerous fine-tuning techniques have been developed. LoRA (Hu et al., 2021) reduces fine-tuning costs via low-rank weight updates, while Textual Inversion (Mokady et al., 2022; Gal et al., 2022) introduces new word embeddings for novel concepts without altering core parameters. DreamBooth (Ruiz et al., 2022) adapts models to specific subjects or styles using a few personalized images, and DreamSync (Sun et al., 2023) enables models to self-improve by learning from their own high-quality outputs. Recently, RLHF (Wang et al., 2025; Gong et al., 2025; Jiang et al., 2025) in Text-to-Vision generation has shown promise as an efficient fine-tuning strategy. In this work, we use several fine-tuning techniques with Generate Any Scene to improve Text-to-Vision generation models.